OpenStudy (anonymous):

Solve for x - 2(x^2 + 1/x) -3 (x-1/x) -4 =0 I remember that when I earlier did this i took something as another variable and then got a quadratic equation from which I solved for the possible values of the taken variable !!! Help me please guys !!!!

OpenStudy (ash2326):

Is this your question \[2(x^2+ \frac{1}{x})-3 (x-\frac{1}{x})-4=0\]

OpenStudy (anonymous):

yes absolutely right

OpenStudy (ash2326):

I think the first term should be \[2 (x^2+\frac{1}{x^2})\] you sure about this?

OpenStudy (anonymous):

oops sorry !! it is x^2 +1/X^2 sorry !! you r correct

OpenStudy (ash2326):

Great :D now Let's solve this

OpenStudy (anonymous):

Anybody Help please !!!!

OpenStudy (ash2326):

we have \[2(x^2+\frac{1}{x^2}) -3(x-\frac{1}{x})-4=0\] we know that \[(x-\frac{1}{x})^2=x^2+\frac{1}{x^2}-2\] so \[(x-\frac{1}{x})^2+2=x^2+\frac{1}{x^2}\] so we have in our equation \[2((x-\frac{1}{x})^2+2) -3(x-\frac{1}{x})-4=0\] Let's substitute \[x-\frac{1}{x}= t\] so we get \[2(t^2+2)-3t-4=0\] we get now \[2t^2+4-3t-4=0\] we get \[2t^-3t=0\] now we get \[t(2t-3)=0\] so either t=0 or t= 3/2 \[x-\frac{1}{x}= 0 or 3/2\] Can you solve now

OpenStudy (anonymous):

Ok great thanks a lot now I remember !!!!!

OpenStudy (ash2326):

we have \[x-\frac{1}{x}=0\] so \[x^2-1=0=> x=\pm 1\] or \[x-1/x=\frac{3}{2}\] so we get \[2x^2-2=3x\] we get \[2x^2-3x-2=0\] so \[2x^2-4x+x-2=0\] \[2x(x-2)+1(x-2)=0\] so \[x=2 or x=-\frac{1}{2}\] so x=1 or -1 or -1/2 or 2

OpenStudy (anonymous):

Ok got it already just needed the first 3 steps !!!! still thanks a lot !!!